# Monogamy of nonlocality from multipartite information causality

### v1

@@ Abstract
The monogamy of nonlocality is one the most intriguing and cryptographically significant predictions of quantum theory. The physical principle of information causality offers a promising means to understand and restrict the extent of nonlocality without invoking the abstract mathematical formalism of quantum theory. In this article, we demonstrate that the original bipartite formulation of information causality cannot imply non-trivial monogamy relations, thereby refuting the previous claims. Nevertheless, we show that the recently proposed multipartite formulation of information causality implies stronger-than-no-signaling monogamy relations. We use these monogamy relations to enhance the security of device-independent quantum key distribution against a no-signaling eavesdropper constrained by information causality.
@@
Introduction:— Quantum theory predicts strong correlations between spatially separated observers, which defy local-causal explanations  [1]. Apart from their foundational significance, nonlocal quantum correlations power several classically inconceivable information processing and cryptographic tasks  [2, 3, 4, 5, 6]. In particular, quantum theory only allows a peculiarly restricted amount of nonlocality [7]. The limits on nonlocality can be estimated using the quantum formalism. However, the abstract mathematical formalism of quantum theory does not offer any insights into the underlying reasons for nature’s restraint on nonlocality. Consequently, numerous efforts have recently been made to recover the set of nonlocal quantum correlations from well-motivated operational principles [8, 9, 10, 11].
@@
The most prominent among such proposals, is the physical principle of information causality ($\mathcal{IC}$) [12], which states that the amount of randomly accessible data cannot exceed the capacity of a classical communication channel even when it is assisted by nonlocal correlations [12, 13, 14]. Thus, $\mathcal{IC}$ forbids stronger-than-quantum nonlocal correlations to a significant extent  [12, 13, 14]. The most outstanding feat of $\mathcal{IC}$ is the recovery of the Tsirelson’s bound on the maximum quantum violation of the Clauser-Horne-Shimony-Holt (CHSH) inequality. While quantum theory satisfies $\mathcal{IC}$, it remains unclear whether all stronger-than-quantum nonlocal correlations necessarily violate $\mathcal{IC}$.
@@
The main challenge in deriving bounds on nonlocal correlations with $\mathcal{IC}$ springs from the inherent dependence of the principle on the communication protocol. A recent contribution [15], made significant headway towards reducing the complexity and broadening the application of $\mathcal{IC}$ by replacing the poorly-scaling concatenation procedure with noisy channels. However, another significant short-coming of the initial formulation $\mathcal{IC}$ is that it is bipartite, and any operational principle which seeks to recover the quantum set of nonlocal correlations must be multipartite [14, 16]. To address this issue, $\mathcal{IC}$ was recently reformulated to apply multipartite communication scenarios and shown to forbid nontrivial stronger-than-quantum multipartite nonlocal correlations [17].
Apart from the extent of nonlocality, a characteristic feature of quantum theory is the monogamy of nonlocality, which limits the distribution of nonlocal correlations among multiple parties. Specifically, monogamy forbids spatially separated parties witnessing strong nonlocal correlations from being strongly correlated with any other party, thereby ensuring the information-theoretic security of Device-Independent Quantum Key Distribution (DIQKD) schemes. This article addresses whether $\mathcal{IC}$ implies quantum-like monogamy relations.
We first show that the original bipartite formulation of $\mathcal{IC}$ cannot imply a non-trivial CHSH monogamy relation beyond no-signaling, refuting previous claims made using this formulation to derive quantum CHSH monogamy from $\mathcal{IC}$  [18, 19, 20]. We then demonstrate that the multipartite formulation of $\mathcal{IC}$ implies stronger than no-signaling CHSH-monogamy relations. In particular, when two parties observe the maximal violation of the CHSH inequality, multipartite $\mathcal{IC}$ forbids any correlation with the third party, thereby retrieving quantum monogamy and guaranteeing information-theoretic security of DIQKD. Whereas for non-maximally nonlocal correlations between two parties, multipartite $\mathcal{IC}$ implies tighter than no-signaling bounds on nonlocal correlations with a third party and bolsters the security of DIQKD against a no-signaling adversary additionally constrained by $\mathcal{IC}$.
@@
Here, $\oplus$ represents the sum modulo 2. The value of the CHSH functional, $\beta(\mathcal{A},\mathcal{B})$, ranges from $1/2$ to $1$, with the maximum local-causal value capped at $3/4$. $\beta(\mathcal{A},\mathcal{B})$ not only detects nonlocal correlations but also quantifies their strength. According to quantum theory, the CHSH inequality can be violated up to $\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})\approx 0.8535$, a characteristic limit known as the Tsirelson’s bound [21]. Correlations which satisfy the no-signaling condition can attain an even higher violation of CHSH inequality up to $\beta_{NS}=1$. In general, correlations which violate the CHSH inequality exhibit several nonclassical features, the most notable and cryptographically significant being their monogamy.
Monogamy of nonlocality restricts the extent to which nonlocal correlations can be shared between multiple spatially separated parties. Specifically, let us consider a tripartite Bell scenario including $\mathcal{A},\mathcal{B}$, the additional party $\mathcal{E}$ with an input $z\in\{0,1\}$ and output $e\in\{0,1\}$. Then monogamy of nonlocality implies that if the two parties $\mathcal{A},\mathcal{B}$ observe nonlocal correlations, such that $\beta(\mathcal{A},\mathcal{B})>3/4$, then the strength of their correlations with $\mathcal{E}$, as measured by the value of the CHSH functional $\beta(\mathcal{B},\mathcal{E})$ (or $\beta(\mathcal{A},\mathcal{E})$) remains limited. This notion is captured by means of a monogamy relation of the generic form,
@@
where $f_{T}^{M}:[1/2,1]\mapsto[0,1]$ is a function specifying the characteristic monogamy relation of given nonlocal theory $T$. Monogamy relations of the form (2) are cryptographically significant as they can be used to derive criteria for ensuring the security of DIQKD against adversaries restricted by the nonlocal theory $T$ [22]. In particular, for the DIQKD protocol based on the CHSH scenario considered in [22, 6], the sufficient condition from ensuring security against individual attacks translates to [23, 24],
@@
$h(p)=-p\log p-(1-p)\log(1-p)$ is the Shonnon’s binary entropy. The condition (3) implies threshold values of $\beta(\mathcal{A},\mathcal{B})$ required for the security, which can be obtained by substituting $f_{T}^{M}$ in (3), with the monogamy relation (2), for any given nonlocal theory $T$.
For instance, correlations satisfying the no-signaling conditions obey the following linear monogamy relation,
@@
Notice that when $\beta(\mathcal{A},\mathcal{B})=\beta_{NS}=1$ then the no-signaling condition implies that $\mathcal{B}$ (and $\mathcal{A}$) must be completely uncorrelated with the third party $\mathcal{E}$, such that, $\beta(\mathcal{B},\mathcal{E})=1/2$. The threshold value of $\beta(\mathcal{A},\mathcal{B})$ for secure DIQKD under with no-signaling monogamy (4) turns out to be $0.881$, which is not realizable with quantum theory [24]. Of particular relevance, quantum theory features a tighter than no signaling, characteristic quadratic monogamy relation [25],
@@
Similar to the no-signaling case, when $\mathcal{A},\mathcal{B}$ observe $\beta(\mathcal{A},\mathcal{B})=\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$, then $\beta(\mathcal{B},\mathcal{E})$ must be $1/2$. In this case, the threshold value of $\beta(\mathcal{A},\mathcal{B})$ for secure DIQKD turns out to be $\approx 0.841$ (3). While (5) was derived for the quantum formalism, we are interested in whether a non-trivial, i.e., tighter than (4), monogamy relation of the form (2) can be derived via a physical principle, without invoking the abstract Hilbert space formalism. Towards this end, we now present the physical principle of information casuality.
Information causality ($\mathcal{IC}$):— The principle of $\mathcal{IC}$ is typically formulated by means of a bipartite communication task, called the $(n\mapsto m)$ random access code (RAC), wherein the parties utilise a nonlocal correlation assisted by a classical communication channel of bounded capacity [26]. Specifically, the sender ($\mathcal{A}$) receives a randomly sampled bit string $\mathbf{x}\equiv\{X_{1},\ldots,X_{n}\}\in\{0,1\}^{n}$ of length $n$. $\mathcal{A}$ then encodes $\mathbf{x}$ onto a classical message $M$ of $m$ bits with $m<n$. The message $M$ is then transmitted through a noisy classical channel with capacity $C\leq m$, such that $\mathcal{B}$ gets $M^{\prime}$. $\mathcal{B}$ then decodes the message $M^{\prime}$ to produce a guess $G_{i}$ about a randomly selected bit of $X_{i}$ of $\mathcal{A}$, where $i\in\{1,\ldots,n\}$. In this set-up, the principle of $\mathcal{IC}$ states that the total potential information $\mathcal{B}$ can gain about the $\mathcal{A}$’s bit string $\mathbf{x}$ cannot exceed the capacity $C$ of the classical communication channel, i.e.,
@@
where $I(X_{i}:G_{i})$ denotes Shannon’s mutual information between $\mathcal{A}^{\prime}s$ $i$-th input $X_{i}$ and the corresponding $\mathcal{B}$’s guess $G_{i}$, $\sum_{i=1}^{n}I(X_{i}:G_{i})$ is the total potentially accessible information, and $C\equiv I(M:M^{\prime})$ is capacity of the noisy classical channel  [^1] We note here that in (6) we are considering the most recent definition proposed in [15], which is a generalization of the original proposal [12]..
Quantum theory satisfies $\mathcal{IC}$, while stronger than quantum nonlocal correlations violate $\mathcal{IC}$, and hence, are ruled out by the principle. Specifically, consider the $(2\mapsto 1)$ RAC, and let $\mathcal{A},\mathcal{B}$ share a no-signaling PR-box correlation [28], defined as: $p(a,b|x,y)=\frac{1}{2}\delta_{a\oplus b,xy}$. The parties then perform the van Dam protocol [4], wherein $\mathcal{A}$ inputs $x=X_{1}\oplus X_{2}$ into her part of the PR-box, and transmits the message $M=a\oplus X_{1}$. Let the classical communication be noiseless such that $C=1$ and $M^{\prime}=M$. $\mathcal{B}$ upon receiving $M$ randomly inputs $y=i-1$ into her PR-box to produce the outcome $G_{i}=M\oplus b$. Since the PR-box satisfies $a\oplus b=xy$, $G_{i}=X_{i}$ and $I(X_{1}:G_{1})=I(X_{2}:G_{2})=C=1$, thereby violating $\mathcal{IC}$ (6). Hence, the PR-box is ruled out by $\mathcal{IC}$. Moreover, any nonlocal correlation $p(a,b|x,y)$ which violates the CHSH inequality beyond the Tsirelson’s bound $\beta(\mathcal{A},\mathcal{B})>\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$ is ruled out by $\mathcal{IC}$, for some $C\in[0,1]$ [12, 15].
Moreover, we can derive an even more general criterion for $\mathcal{IC}$ for the causal structure associated with the $(n\mapsto m)$ RAC using the technique described in [29] to compute $\mathcal{IC}$-like222satisfied by both classical and quantum theory. information theoretic constraints for arbitrary causal structures. Specifically, the method [29] returns the following generalized criterion for $\mathcal{IC}$ which takes into non-uniform priors and arbitrary decoding protocols,
@@
Figure 1: The graphic depicts the causal structure represented as a Directed Acyclic Graph (DAG) associated with the communication task for the multipartite $\mathcal{IC}$ criterion (8), entailing $N-1$ senders and a receiver. The parties have access to a pre-shared entangled quantum state $\rho$ (green square). The senders $\{\mathcal{A}_{k}\}^{N-1}_{k=1}$ receive inputs $\{\{X^{k}_{j}\}^{n}_{j=1}\}^{N-1}_{k=1}$ (blue disks), and transmit messages $\{M_{k}\}^{N-1}_{k=1}$ (pink disks) through binary-symmetric noisy classical channels with parameters $\{\epsilon_{k}\}^{N}_{k=1}$, to the receiver $\mathcal{B}$, respectively. Upon receiving the $N-1$ potentially noisy messages $\{M^{\prime}_{k}\}^{N-1}_{k=1}$, the receiver computes guess $G_{j}$ (purple disk) about a joint function $f_{j}(\{X^{k}_{j}\}^{N-1}_{k=1})$, based on a randomly select input $j\in\{1,\ldots n\}$ (green disk).
Multipartite information causality:— Consider a communication task entailing $N$ spatially separated parties wherein $N-1$ senders $\{\mathcal{A}_{k}\}^{N-1}_{k=1}$ transmit information to a receiver $\mathcal{B}$. Similar to the $(n\mapsto m)$ RAC, each sender $\mathcal{A}_{k}$ receives an $n$-bit string $\mathbf{x}^{k}=(X_{1}^{k},X_{2}^{k},\cdots,X_{n}^{k})$ as input which is encoded on to a $m<n$-bit classical message $M_{k}$ and transmitted to $\mathcal{B}$ via a potentially noisy communication channel with capacity $C_{k}<m$. $\mathcal{B}$ in turn gets the noisy messages $\{M_{k}^{\prime}\}$, and based on randomly chosen input $j\in\{1,\ldots,n\}$ he produces a guess $G_{j}$ about a function of the form $f_{j}(X_{j}^{1},X_{j}^{2},\cdots,X_{j}^{N-1})$. The directed acyclic graph representing the causal structure associated with this communication task in presented in Figure 1. Observe that in contrast to bipartite $(n\mapsto m)$ RAC wherein $\mathcal{B}$ attempts to guess a randomly chosen bit of $\mathcal{A}$, this task is multipartite, as $\mathcal{B}$ guesses a joint function $f_{j}$ of the $j$th input bits of the $N-1$ senders $\{\mathcal{A}_{k}\}^{N-1}_{k=1}$. Consequently, the bipartite criteria of $\mathcal{IC}$ presented above (6) and (7) fail to discard post-quantum correlations in this set-up. Instead, let us consider the following multipartite criterion for $\mathcal{IC}$,
@@
where $\mathbf{M^{\prime}}=(M_{1}^{\prime},M_{2}^{\prime},\cdots,M_{N-1}^{\prime})$ denotes the tuple of the messages reaching the receiver through the $N$ potentially noisy classical channels. Note that, the criterion (8) generalizes the once presented in [17] by allowing for noisy classical communication. In the supplementary material, we show that it is satisfied by theories, including classical and quantum theory, which satisfy the information theoretic axioms of $\mathcal{IC}$.
In contrast to bipartite criteria (6), (7), the multipartite criterion (8) for $\mathcal{IC}$ forbids stronger than quantum no-signaling correlations in this set-up (Figure 1). Specifically, let us consider the simplest tripartite version of the multipartite communication task presented above, with $N=3$ and $n=2$. Let the two senders $\mathcal{A},\mathcal{E}$ and the receiver $\mathcal{B}$ share a tripartite no-signaling correlation $p(a,b,e|x,y,z)$ satisfying $a\oplus b\oplus e=xy\oplus zy$. $\mathcal{A},\mathcal{E}$ input uncorrelated bits $x=X_{0}^{1}\oplus X_{1}^{1}$, $z=X_{0}^{2}\oplus X_{1}^{2}$, and transmit the message $M_{1}=a\oplus x_{0}^{1}$ and $M_{2}=e\oplus x_{0}^{2}$ via noiseless communication channels such that $C_{1}=C_{2}=1$, respectively. Upon receiving the messages $\{M_{k}\}^{2}_{k=1}$, $\mathcal{B}$ then inputs $y=j-1$ and produces the guess $G_{j}=M_{1}\oplus M_{2}\oplus b$. Consequently, we have that $I(X_{i}^{k}:X_{i}^{3-k},M_{1},M_{2},G_{i})=C_{k}=1$, and $I(X_{i+1}^{k},\dots,X_{n}^{k}:X_{i}^{k})=0$ for all $k,i\in\{1,2\}$, which violates the multipartite $\mathcal{IC}$ criterion (8), thereby, discarding the no-signaling correlation. We are now prepared to test whether non-trivial monogamy relations of the form (2) can be derived from the bipartite (6),(7) and multiparty (8) criteria for $\mathcal{IC}$.
Optimal slice:— To retrieve monogamy relations of the form (2), we need to find the maximum value of $\beta(\mathcal{B},\mathcal{E})$ given $\beta(\mathcal{A},\mathcal{B})$ over all tripartite no-signaling correlations which satisfy the respective non-linear and protocol-dependent $\mathcal{IC}$ criteria. Since the convex polytope of tripartite no-signaling correlations has $53856$ extremal points, these optimization problems are particularly complex. We now present a useful Lemma which significantly reduces this complexity and allows us to restrict to a two-parameter slice of the tripartite no-signaling polytope,

### v2

# Security of device-independent quantum key distribution via monogamy relations from multipartite information causality

@@ Abstract
The information causality principle ($\mathcal{IC}$) was originally proposed as a promising means to understand the extent of quantum nonlocality without invoking the full mathematical formalism of quantum theory. Beyond the foundational significance, the problem of bounding nonlocal correlations by reasonable physical principles has meaningful practical consequences, particularly for device-independent (DI) cryptographic security. In this work, we advance in this direction, demonstrating that the $\mathcal{IC}$ is enough to ensure DI security on quantum key distribution (QKD) protocols. Security is proven for a range of theoretically quantum-attainable parameters against individual attacks by a potentially post-quantum eavesdropper. This result follows as a consequence of a strong form of monogamy of Bell’s inequality violations, which we have proven to be implied by the recently proposed multipartite formulation for $\mathcal{IC}$. Additionally, we demonstrated that the original bipartite formulation of $\mathcal{IC}$ fails to imply monogamy relations and hence, ensure security of DIQKD, thus stressing the necessity of the multipartite framework.
@@
Introduction:— Quantum theory predicts strong correlations between spatially separated observers, which defy local-causal explanations [1]. Apart from their foundational significance, nonlocal quantum correlations power several classically inconceivable information processing tasks [2, 3, 4, 5, 6, 7]. In particular, quantum theory only allows a peculiarly restricted amount of nonlocality [8], which can be estimated using the quantum formalism. However, the abstract mathematical formalism of quantum theory does not offer any insights into the underlying reasons for nature’s restraint on nonlocality. Consequently, numerous efforts have recently been made to recover the set of nonlocal quantum correlations from well-motivated physical principles [9, 10, 11, 12].
@@
The most prominent among such proposals, is the physical principle of information causality ($\mathcal{IC}$) [13], which states that the amount of randomly accessible data cannot exceed the capacity, $C$, of a classical communication channel even when it is assisted by nonlocal correlations [13, 14, 15]. As originally suggested in 2009 [13], $\mathcal{IC}$ extends the idea of NS principle to communicating scenarios – whereas $\rightarrow 0, 𝒞\mathcal{IC}$ reduces to no-signalling. Thus, $\mathcal{IC}$ forbids stronger-than-quantum nonlocal correlations to a significant extent [13, 14, 15]. The most outstanding feat of $\mathcal{IC}$ is the recovery of the Tsirelson’s bound on the maximum quantum violation of the Clauser-Horne-Shimony-Holt (CHSH) inequality. While quantum theory satisfies $\mathcal{IC}$, it remains unclear whether all stronger-than-quantum nonlocal correlations necessarily violate $\mathcal{IC}$.
@@
The main challenge in deriving bounds on nonlocal correlations with $\mathcal{IC}$ springs from the inherent dependence of the principle on the communication protocol. A recent contribution [16], made significant headway towards reducing the complexity and broadening the application of $\mathcal{IC}$ by replacing the poorly-scaling concatenation procedure with noisy channels. However, another significant short-coming of the initial formulation $\mathcal{IC}$ is that it is bipartite, and any operational principle which seeks to recover the quantum set of nonlocal correlations must be multipartite [15, 17]. To address this issue, $\mathcal{IC}$ was recently reformulated to apply to multipartite communication scenarios and shown to forbid nontrivial stronger-than-quantum multipartite nonlocal correlations [18].
Although the precise boundary of quantum nonlocality remains unexplained in terms of fundamental physical principles, such as information causality ($\mathcal{IC}$), its deep connection with cryptographic tasks constitutes one of the most striking practical consequences of quantum theory [6, 19]. In particular, the quantum key distribution (QKD) framework ensures security based solely on the laws of quantum mechanics without relying on any assumptions about the computational power of a potential eavesdropper. The device-independent quantum key distribution (DIQKD) framework extends this idea further by making no assumptions about the physical description of the devices used by the parties, instead grounding security in fundamental operational principles [6]. Under this paradigm, when parties are constrained by the no-signaling principle—prohibiting communication faster than the speed of light—security proofs may remain valid even in the face of a hypothetical post-quantum breakthrough that grants an adversary access to supra-quantum resources. On the other hand, a broad class of supra-quantum no-signaling correlations has been consistently shown to be likely non-physical, as they would lead to implausible consequences —for instance, making the communication complexity problem trivial[4] or even violating the $\mathcal{IC}$ principle [13]. Consequently, this fact raises the fundamental question of whether device-independent security proofs can be established based on alternative physical principles.
In this work, we take a significant step toward demonstrating that the $\mathcal{IC}$ principle ensures the security of DIQKD protocols. In particular, by employing a refined multipartite informational-theoretic approach, we have shown that the so-called CHSH protocol [6] is safe when parties respect the multipartite operational form for the $\mathcal{IC}$ principle. Notably, the security holds for a range of quantum-attainable violations of the CHSH inequality. This result follows directly from a nontrivial monogamy relation for CHSH inequalities, which we derive from the multipartite formulation of $\mathcal{IC}$. In particular, when two parties observe the maximal violation of the CHSH inequality, multipartite $\mathcal{IC}$ strictly prohibits any correlation with a third party, thereby fully recovering the quantum monogamy of Bell inequality violations. For non-maximal violations, multipartite $\mathcal{IC}$ imposes tighter than no-signaling bounds, effectively limiting nonlocal correlations with a third party and enhancing the security of DIQKD against an adversary constrained by $\mathcal{IC}$. Additionally, we demonstrate that previous bipartite formulations of $\mathcal{IC}$ fail to impose the same monogamy constraints on CHSH inequalities, highlighting the necessity of the multipartite framework for $\mathcal{IC}$.
@@
Here, $\oplus$ represents the sum modulo 2. The value of the CHSH functional, $\beta(\mathcal{A},\mathcal{B})$, not only determines the presence of nonlocal correlations but also quantifies their strength. Specifically, the completely uncorrelated case yields $1/2$, whereas the maximum value allowed by local causality is capped at $3/4$. According to quantum theory, the CHSH inequality can be violated up to $\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})\approx 0.8535$, a characteristic limit known as the Tsirelson’s bound [20]. Correlations which satisfy the no-signaling condition can attain an even higher violation of CHSH inequality up to $\beta_{NS}=1$. In general, correlations which violate the CHSH inequality exhibit several nonclassical features, the most notable and cryptographically significant being their monogamy.
Monogamy of nonlocality restricts the extent to which nonlocal correlations can be shared between multiple spatially separated parties. Specifically, let us consider a tripartite Bell scenario including $\mathcal{A},\mathcal{B}$, and an additional party $\mathcal{E}$ with an input $z\in\{0,1\}$ and an output $e\in\{0,1\}$. Then monogamy of nonlocality implies that if the two parties $\mathcal{A},\mathcal{B}$ observe nonlocal correlations, such that $\beta(\mathcal{A},\mathcal{B})>3/4$, then the strength of their correlations with $\mathcal{E}$, as measured by the value of the CHSH functional $\beta(\mathcal{B},\mathcal{E})$ (or $\beta(\mathcal{A},\mathcal{E})$) remains limited. This notion is captured by means of a monogamy relation of the generic form,
@@
where $f_{T}^{M}:[1/2,1]\mapsto[0,1]$ is a function specifying the characteristic monogamy relation of given nonlocal theory $T$. Monogamy relations of the form (2) are cryptographically significant as they can be used to derive criteria for ensuring the security of DIQKD against adversaries restricted by the nonlocal theory $T$ [21]. In particular, for the DIQKD protocol based on the CHSH set-up considered in [21, 6], the sufficient condition from ensuring security against individual attacks translates to [22, 23],
@@
where $h(p)=-p\log p-(1-p)\log(1-p)$ is the Shannon’s binary entropy. The condition (3) implies threshold values of $\beta(\mathcal{A},\mathcal{B})$ required for the security, which can be obtained by substituting $f_{T}^{M}$ in (3), with the monogamy relation (2), for any given nonlocal theory $T$.
For instance, correlations satisfying the no-signaling conditions obey the following linear monogamy relation[24],
@@
Notice that when $\beta(\mathcal{A},\mathcal{B})=\beta_{NS}=1$, the no-signaling condition implies that $\mathcal{B}$ (and $\mathcal{A}$) must be completely uncorrelated with the third party $\mathcal{E}$, such that, $\beta(\mathcal{B},\mathcal{E})=1/2$. The threshold value of $\beta(\mathcal{A},\mathcal{B})$ for secure DIQKD with no-signaling monogamy (4) turns out to be $\approx 0.881$, which is not realizable with quantum theory [23]. Of particular relevance, quantum theory features a tighter than no signaling, characteristic quadratic monogamy relation [25],
@@
Similar to the no-signaling case, when $\mathcal{A},\mathcal{B}$ observe $\beta(\mathcal{A},\mathcal{B})=\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$, then $\beta(\mathcal{B},\mathcal{E})$ must be $1/2$. In this case, the threshold value of $\beta(\mathcal{A},\mathcal{B})$ for secure DIQKD turns out to be $\approx 0.841$ (3). While (5) holds for quantum theory, we are interested in whether a non-trivial, i.e., tighter than (4), monogamy relation of the form (2) can be derived via a physical principle, without invoking the abstract Hilbert space formalism. Towards this end, we now present the physical principle of information causality.
Information causality ($\mathcal{IC}$):— The principle of $\mathcal{IC}$ is typically formulated by means of a bipartite communication task, namely, the $(n\mapsto m)$ random access code (RAC), wherein the parties utilise a nonlocal correlation assisted by a classical communication channel of bounded capacity [26]. Specifically, the sender ($\mathcal{A}$) receives a randomly sampled bit string $\mathbf{x}\equiv(X_{1},\ldots,X_{n})\in\{0,1\}^{n}$ of length $n$. $\mathcal{A}$ then encodes $\mathbf{x}$ onto a classical message $M$ of $m$ bits with $m<n$. The message $M$ is then transmitted through a noisy classical channel with capacity $C\leq m$, such that $\mathcal{B}$ gets $M^{\prime}$. $\mathcal{B}$ then decodes the message $M^{\prime}$ to produce a guess $G_{i}$ about a randomly selected bit of $X_{i}$ of $\mathcal{A}$, where $i\in\{1,\ldots,n\}$. In this set-up, the principle of $\mathcal{IC}$ states that the total potential information $\mathcal{B}$ can gain about the $\mathcal{A}$’s bit string $\mathbf{x}$ cannot exceed the capacity $C$ of the classical communication channel, i.e.,
@@
where $I(X_{i}:G_{i})$ denotes Shannon’s mutual information between $\mathcal{A}$’s $i$-th input $X_{i}$ and the corresponding $\mathcal{B}$’s guess $G_{i}$, $\sum_{i=1}^{n}I(X_{i}:G_{i})$ is the total potentially accessible information.  [^1] We note here that in (6) we are considering the most recent definition proposed in [16], which is a generalization of the original proposal [13]..
Quantum theory satisfies $\mathcal{IC}$, while stronger than quantum nonlocal correlations violate $\mathcal{IC}$, and hence, are ruled out by the principle. Specifically, let us consider the $(2\mapsto 1)$ RAC, and let $\mathcal{A},\mathcal{B}$ share a no-signaling PR-box correlation [28], defined as: $p(a,b|x,y)=\frac{1}{2}\delta_{a\oplus b,xy}$. The parties then perform the van Dam protocol [4], wherein $\mathcal{A}$ inputs $x=X_{1}\oplus X_{2}$ into her part of the PR-box, and transmits the message $M=a\oplus X_{1}$. Let the classical communication be noiseless such that $C=1$ and $M^{\prime}=M$. $\mathcal{B}$ upon receiving $M$ randomly inputs y=i1y=i-1 into his PR-box to produce the outcome $G_{i}=M\oplus b$. Since the PR-box satisfies $a\oplus b=xy$, $G_{i}=X_{i}$ and $I(X_{1}:G_{1})=I(X_{2}:G_{2})=C=1$, thereby violating the $\mathcal{IC}$ criterion (6). Hence, the PR-box is ruled out by $\mathcal{IC}$. Moreover, any nonlocal correlation $p(a,b|x,y)$ which violates the CHSH inequality beyond the Tsirelson’s bound $\beta(\mathcal{A},\mathcal{B})>\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$ is ruled out by $\mathcal{IC}$, for some $C\in[0,1]$ [13, 16].
Moreover, we can derive an even more general criterion for $\mathcal{IC}$ by considering the causal structure associated with the $(n\mapsto m)$ RAC and using the technique described in [29] to compute $\mathcal{IC}$-like 222satisfied by both classical and quantum theory. information theoretic constraints for arbitrary causal structures. Specifically, the method [29] returns the following generalized criterion for $\mathcal{IC}$ which takes into non-uniform priors and arbitrary decoding protocols,
@@
Figure 1: The graphic depicts the causal structure represented as a Directed Acyclic Graph (DAG) associated with the communication task for the multipartite $\mathcal{IC}$ criterion (8), entailing N1N-1 senders and a receiver. The parties have access to a pre-shared entangled quantum state $\rho$ (green square). The senders $\{\mathcal{A}_{k}\}^{N-1}_{k=1}$ receive inputs $\{\{X^{k}_{j}\}^{n}_{j=1}\}^{N-1}_{k=1}$ (blue disks), and transmit messages $\{M_{k}\}^{N-1}_{k=1}$ (pink disks) through binary-symmetric noisy classical channels with parameters $\{\epsilon_{k}\}^{N}_{k=1}$, to the receiver $\mathcal{B}$, respectively. Upon receiving the N1N-1 potentially noisy messages $\{M^{\prime}_{k}\}^{N-1}_{k=1}$, the receiver computes guess $G_{j}$ (purple disk) about a joint function $f_{j}(\{X^{k}_{j}\}^{N-1}_{k=1})$, based on a randomly selected input $j\in\{1,\ldots n\}$ (green disk).
Multipartite information causality:— Consider a communication task entailing $N$ spatially separated parties wherein N1N-1 senders $\{\mathcal{A}_{k}\}^{N-1}_{k=1}$ transmit information to a receiver $\mathcal{B}$. Similar to the $(n\mapsto m)$ RAC, each sender $\mathcal{A}_{k}$ receives an $n$-bit string $\mathbf{x}^{k}=(X_{1}^{k},X_{2}^{k},\cdots,X_{n}^{k})$ as input which is encoded on to a $m<n$-bit classical message $M_{k}$ and transmitted to $\mathcal{B}$ via a potentially noisy communication channel with capacity $C_{k}<m$. $\mathcal{B}$ in turn gets the noisy messages $\{M_{k}^{\prime}\}$, and based on randomly chosen input $j\in\{1,\ldots,n\}$ he produces a guess $G_{j}$ about a function of the form $f_{j}(X_{j}^{1},X_{j}^{2},\cdots,X_{j}^{N-1})$. The directed acyclic graph representing the causal structure associated with this communication task in presented in Figure 1. Observe that in contrast to bipartite $(n\mapsto m)$ RAC wherein $\mathcal{B}$ attempts to guess a randomly chosen bit of $\mathcal{A}$, this task is multipartite, as $\mathcal{B}$ guesses a joint function $f_{j}$ of the $j$th input bits $\{X_{j}^{k}\}^{N-1}_{k=1}$ of the N1N-1 senders $\{\mathcal{A}_{k}\}^{N-1}_{k=1}$. Consequently, the bipartite criteria of $\mathcal{IC}$ presented above (6) and (7) fail to discard post-quantum correlations in this set-up. Instead, let us consider the following multipartite criterion for $\mathcal{IC}$,
@@
where $\mathbf{M^{\prime}}=(M_{1}^{\prime},M_{2}^{\prime},\cdots,M_{N-1}^{\prime})$ denotes the tuple of the messages reaching the receiver through the $N$ potentially noisy classical channels. Note that the criterion (8) generalizes the one presented in [18] by accounting for the noisy received message on the receiver’s knowledge about the initial data. In the supplementary material, we show that theories—such as classical and quantum mechanics—that satisfy the information-theoretic axioms of $\mathcal{IC}$ also fulfill this condition and further explain why this refinement is essential for deriving monogamy relations.
In contrast to bipartite criteria (6), (7), the multipartite criterion (8) for $\mathcal{IC}$ forbids stronger than quantum no-signaling correlations in this set-up (Figure 1). Specifically, let us consider the simplest tripartite version of the multipartite communication task presented above, with $N=3$ and $n=2$. Let the two senders $\mathcal{A},\mathcal{E}$ and the receiver $\mathcal{B}$ share a tripartite no-signaling correlation $p(a,b,e|x,y,z)$ satisfying $a\oplus b\oplus e=xy\oplus zy$. $\mathcal{A},\mathcal{E}$ input uncorrelated bits $x=X_{0}^{1}\oplus X_{1}^{1}$, $z=X_{0}^{2}\oplus X_{1}^{2}$, and transmit the message $M_{1}=a\oplus W_{0}^{1}$ and $M_{2}=e\oplus X_{0}^{2}$ via noiseless communication channels such that $C_{1}=C_{2}=1$, respectively. Upon receiving the messages $\{M_{k}\}^{2}_{k=1}$, $\mathcal{B}$ then inputs y=j1y=j-1 and produces the guess $G_{j}=M_{1}\oplus M_{2}\oplus b$. Consequently, we have that $I(X_{i}^{k}:X_{i}^{3-k},M_{1},M_{2},G_{i})=C_{k}=1$, and $I(X_{i+1}^{k},\dots,X_{n}^{k}:X_{i}^{k})=0$ for all $k,i\in\{1,2\}$, which violates the multipartite $\mathcal{IC}$ criterion (8), thereby, ruling out the no-signaling correlation. We are now ready to examine the secrecy of cryptographic protocols when parties are limited by the $\mathcal{IC}$ principle.
Optimal slice:— As introduced in the preamble, the secrecy of DIQKD protocols has close relations with monogamy of Bell inequalities violation, as follows from (3). Consequently, to assess the $\mathcal{IC}$ principle’s ability to ensure security, we investigate its connection with monogamy relations in the form (2). In this regard, the problem reads as finding the maximum value of $\beta(\mathcal{B},\mathcal{E})$ given $\beta(\mathcal{A},\mathcal{B})$ over all tripartite no-signaling correlations which satisfy the respective non-linear and protocol-dependent $\mathcal{IC}$ criteria. Thus, such procedure shall be performed for each bipartite (6),(7) and multipartite (8) criteria. Since the convex polytope of tripartite no-signaling correlations has $53856$ extremal points [31], these optimization problems are particularly complex. We now present a useful Lemma that significantly reduces this complexity and allows us to restrict to a two-parameter slice of the tripartite no-signaling polytope,